1/3x + 6/5 = 5x - 2/5
Answer by checkley71(8403)

1/3x+6/5=5x-2/5
1/3x-5x=-2/5-6/5
(x-15x)/3=-8/5
-14x/3=-8/5 now cross multiply:
-14x*5*=-24
-70x=-24
x=-24/-70
x=12/35 answer.